Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Carlo Migoni at Sun Feb 24 13:00:30 MET 2002) Success

Effelsberg (A. Kraus at Mon Feb 25 08:53:45 MET 2002) Minor failures
We had to stop repeatedly due to snowfall: UT0228-0709, UT0830-0914, and UT0944-1030.

Torun (K.Borkowski at Mon Feb 25 09:28:43 MET 2002) Minor failures
BBC7 out of order. No other known problems.

Westerbork (Tony Foley at Mon Feb 25 14:12:35 MET 2002) Minor failures
Some snow during this observation. Several scans which were outside our hour angle range were replaced with local calibrators. Strong EMI about 04:23 UT.

Jodrell_Bank (Alastair Gunn at Tue Feb 26 14:36:12 MET 2002) Minor failures
Telescope offline at start of experiment due to high winds but schedule was running. Back online at about 23:35 UT. No other known problems.

Cambridge (Alastair Gunn at Tue Feb 26 14:36:36 MET 2002) Failure
Cambridge did not observe due to serious fault with recorder.

Westerbork (Tony Foley at Fri Mar 1 11:52:52 MET 2002) Minor failures
Some scans out of hour angle range. replaced with local calibrators